OpenStudy (anonymous):

The right triangular flag of a sports club was designed to have a base length of 4 ft and height of 6 ft. For a sports event, the club made a new flag by doubling the base and height of the flag. The area of the new flag is ______ times larger than the original flag.

OpenStudy (btaylor):

4 times larger. original: (.5 x 4 x 6) = 12 new: (.5 x 8 x 12) = 48

Parth (parthkohli):

\( \color{Black}{\Rightarrow 0.5 \times 4 \times 6 = ? }\) \( \color{Black}{\Rightarrow 0.5 \times 8 \times 12 = ?}\)

OpenStudy (anonymous):

yep thats wat i got! thanks

Parth (parthkohli):

Then, you have to find \(48 \div 12\) to get the final answer.

OpenStudy (anonymous):

great thanks ur awesome!